# Complexity Bounds for Smooth Convex Multiobjective Optimization

### v1

@@ 4.2.1 Oblivious One-Step Gradient Methods: An $\Omega(1/T)$ Min-Iterate Bound
Step 1: Spectral hard instance (diagonal is sufficient). By Lemma 3.13, for an oblivious one-step method on a quadratic with spectrum in $[0,L]$ we have $\nabla g(x^{(t)})=H\,p_{t}(H)\,(x^{(0)}-x^{\star})$ with $p_{t}(\zeta)=\prod_{k=0}^{t-1}(1-\alpha_{k}\zeta)$ and $\alpha_{k}\leq 1/L$. For any fixed stepsizes, the worst case over instances is obtained by placing an eigenvalue at a maximizer $\zeta^{\star}\in[0,L]$ of $\zeta\big{|}\prod_{k=0}^{t-1}(1-\alpha_{k}\zeta)\big{|}$ and aligning $x^{(0)}-x^{\star}$ with the corresponding eigenvector; in particular, a diagonal (even one-dimensional) $H=\zeta^{\star}$ suffices. Thus the lower bound reduces to the product-form extremal
$\inf_{\{\alpha_{k}\}_{k=0}^{t-1}\subset[0,1/L]}\ \max_{\zeta\in[0,L]}\ \zeta\ \Bigl{|}\prod_{k=0}^{t-1}(1-\alpha_{k}\zeta)\Bigr{|}\ \geq\ \frac{L}{4(t+1)},$
proved in Appendix A.1. Taking the minimum over $t\in\{0,\dots,T\}$ is achieved at $t=T$.
Step 2: Lifting to MOO. Using the construction of Theorem 4.5 with this $L$-smooth convex quadratic $g$, we obtain the desired MOO lower bound. ∎
4.2.2 Fully Adaptive Methods: A Universal Last-Iterate Lower Bound and an Open Problem
When the algorithm can be fully adaptive, the error polynomial $p_{t}$ no longer has the restrictive product form. This leads to a weaker, but more general, lower bound based on Markov’s inequality for polynomials.
Theorem 4.4 (Universal Convex Lower Bound).
Fix $L,R>0$ and $T\in\mathbb{N}$. For any span first-order method run for $T$ steps, there exists an MOO instance with $L$-smooth convex objectives and $\mathrm{dist}(x^{(0)},\mathcal{P})=R$ such that the last iterate obeys
@@ 4.2.2 Fully Adaptive Methods: A Universal Last-Iterate Lower Bound and an Open Problem
Step 1: The Hard Scalar Instance. We use the same hard quadratic $g(x)$ as in the proof of Theorem 4.3. For a general span method, the error polynomial $p_{t}$ from Lemma 3.13 can be any polynomial of degree at most $t$ with $p_{t}(0)=1$. The lower bound comes from the extremal problem:
This result is a direct consequence of the Markov brothers’ inequality, which bounds the derivative of a polynomial. The full derivation is provided in Appendix A.2. The minimum over $t\in\{0,\dots,T\}$ is achieved at $t=T$, yielding $\frac{LR}{2(T+1)^{2}}$.
Step 2: Lifting to MOO. The result follows by applying the construction of Theorem 4.5, which translates the scalar gradient norm lower bound into a lower bound on the Pareto stationarity gap. ∎
@@
This $\Omega(1/T^{2})$ bound is consistent with (but weaker than) the $\mathcal{O}(LR/\sqrt{T})$ gradient-norm upper bound of vanilla gradient descent, and leaves a gap to the $\mathcal{O}(LR/T)$ upper bound achieved by AGD.
Open Problem. Does a universal $\Omega(1/T)$ lower bound hold for the Pareto stationarity gap $\min_{t\leq T}\mathcal{G}(x^{(t)})$ for all span first-order methods in smooth convex MOO, or can an adaptive algorithm achieve a rate faster than $\mathcal{O}(1/T)$?
4.3 A Construction for Non-Degenerate Instances
A potential criticism of lower bounds is that they might only hold on degenerate problem instances (e.g., where all $f_{i}$ are identical, or the Pareto set is a single point). We now show that our bounds hold for robust, non-degenerate problems.
Theorem 4.5 (Robustness of Lower Bounds).
The lower bounds stated in Theorems 4.1, 4.3, and 4.4 hold for MOO instances where:
@@ 4.3 A Construction for Non-Degenerate Instances
    The objective functions $\{f_{i}\}_{i=1}^{m}$ are distinct.
@@
    The Pareto set $\mathcal{P}$ is a non-singleton convex polytope.
@@
Let a span first-order method and $T$ be given. We construct a hard MOO instance. Decompose the space as $\mathbb{R}^{d}=V\oplus W$, where $\dim(V)\geq T+1$ and $\dim(W)\geq m-1$. Any $x\in\mathbb{R}^{d}$ is written as $x=x_{V}+x_{W}$. Let $g:V\to\mathbb{R}$ be the hard scalar quadratic function from the previous proofs (either smooth convex or $\mu$-strongly convex as needed). In $W$, choose $m$ affinely independent points $\{a_{1},\dots,a_{m}\}$.
We define the objectives $f_{i}:\mathbb{R}^{d}\to\mathbb{R}$ for $i=1,\dots,m$ as:
$f_{i}(x)\;=\;g(x_{V})\;+\;\frac{\gamma}{2}\,\|x_{W}-a_{i}\|^{2}.$
The parameter $\gamma>0$ is chosen to ensure the desired smoothness and strong convexity.
  • Strongly convex case: Let $g$ be $L$-smooth and $\mu$-strongly convex. We set $\gamma=\mu$. The Hessian is $\nabla^{2}f_{i}(x)=\mathrm{diag}\bigl{(}\nabla^{2}g(x_{V}),\mu I_{W}\bigr{)}$. Since $\left\lVert\nabla^{2}g\right\rVert_{\mathrm{op}}\leq L$ and its eigenvalues are at least $\mu$, each $f_{i}$ is $L$-smooth and $\mu$-strongly convex.
  • Convex case: Let $g$ be $L$-smooth and convex. We set $\gamma=L$. The Hessian is $\nabla^{2}f_{i}(x)=\mathrm{diag}\bigl{(}\nabla^{2}g(x_{V}),LI_{W}\bigr{)}$. Since $\left\lVert\nabla^{2}g\right\rVert_{\mathrm{op}}\leq L$, each $f_{i}$ is $L$-smooth and convex.
The gradient is $\nabla f_{i}(x)=(\nabla g(x_{V}),\gamma(x_{W}-a_{i}))$. A convex combination of gradients is $\sum_{i}\lambda_{i}\nabla f_{i}(x)=\left(\nabla g(x_{V}),\gamma\left(x_{W}-\sum_{i}\lambda_{i}a_{i}\right)\right)$. The squared Pareto stationarity gap is therefore
$\mathcal{G}(x)^{2}\;=\;\min_{\lambda\in\Delta^{m}}\left\|\sum_{i}\lambda_{i}\nabla f_{i}(x)\right\|^{2}=\|\nabla g(x_{V})\|^{2}\;+\;\gamma^{2}\min_{\lambda\in\Delta^{m}}\left\|x_{W}-\sum_{i=1}^{m}\lambda_{i}a_{i}\right\|^{2}.$
This implies that $\mathcal{G}(x)\geq\|\nabla g(x_{V})\|$. A point $x$ is Pareto stationary ($\mathcal{G}(x)=0$) if and only if $\nabla g(x_{V})=0$ and $x_{W}\in\mathrm{conv}\{a_{1},\dots,a_{m}\}$. The Pareto set is thus $\mathcal{P}=\{x_{V}^{\star}\}\times\mathrm{conv}\{a_{i}\}$, which is a non-singleton polytope.
We set the initial point to $x^{(0)}=x_{V}^{(0)}+a_{1}$, where $x_{V}^{(0)}$ is the starting point for the scalar hard instance, satisfying $\|x_{V}^{(0)}-x_{V}^{\star}\|=R$. The distance to the Pareto set is $\mathrm{dist}(x^{(0)},\mathcal{P})=\min_{p\in\mathcal{P}}\|x^{(0)}-p\|=\|(x_{V}^{(0)}+a_{1})-(x_{V}^{\star}+a_{1})\|=\|x_{V}^{(0)}-x_{V}^{\star}\|=R$.
Since the $V$-component of any gradient $\nabla f_{i}(x^{(j)})$ is always $\nabla g(x_{V}^{(j)})$, the update for $x_{V}$ is independent of $x_{W}$ and is identical to that of the algorithm run on the scalar problem $g$. Thus, the sequence $\{x_{V}^{(t)}\}$ is the hard sequence from the scalar lower bounds. Since $\mathcal{G}(x^{(t)})\geq\left\lVert\nabla g(x_{V}^{(t)})\right\rVert$, the scalar lower bounds apply directly to the stationarity gap of this non-degenerate MOO problem. ∎

### v2

@@ 4.4.1 Oblivious One-Step Gradient Methods: An $\Omega(1/T)$ Min-Iterate Bound
By Lemma 4.2(b), for the given stepsize schedule there exists an $L$-smooth convex quadratic $g$ and an initialization $u^{(0)}$ such that
$\min_{0\leq t\leq T}\|\nabla g(u^{(t)})\|\;\geq\;\frac{LR}{4(T+1)}\qquad\text{and}\qquad\mathrm{dist}\!\bigl(u^{(0)},\arg\min g\bigr)=R.$
Choose $u^{\star}\in\arg\min g$ to be a closest minimizer to $u^{(0)}$, so that $\|u^{(0)}-u^{\star}\|=\mathrm{dist}(u^{(0)},\arg\min g)=R$.
Apply Theorem 4.1 with $\gamma=L$ to lift $g$ into a non-degenerate MOO instance $(f_{1},\dots,f_{m})$. By Theorem 4.1(iv), the (weak) Pareto set satisfies
$\mathcal{P}=(\arg\min g)\times\mathrm{conv}\{a_{1},\dots,a_{m}\}.$
Pick $w^{(0)}\in\mathrm{conv}\{a_{1},\dots,a_{m}\}$ (e.g., $w^{(0)}=a_{1}$) and set $x^{(0)}=(u^{(0)},w^{(0)})\in V\oplus W$. Then
$\mathrm{dist}\!\bigl(x^{(0)},\mathcal{P}\bigr)=\mathrm{dist}\!\bigl(u^{(0)},\arg\min g\bigr)=R.$
Finally, Theorem 4.1(iii) gives $\mathcal{G}(x)\geq\|\nabla g(x_{V})\|$ for all $x$. Hence,
$\min_{0\leq t\leq T}\mathcal{G}(x^{(t)})\;\geq\;\min_{0\leq t\leq T}\|\nabla g(x_{V}^{(t)})\|\;\geq\;\frac{LR}{4(T+1)},$
as claimed. ∎
4.4.2 Oblivious span methods: polynomial-degree vs. geometric lower bounds
For oblivious span methods, the error polynomial $p_{t}$ is no longer restricted to have the product form of the one-step class, but it remains instance-independent. This leads to a universal $1/T^{2}$ last-iterate lower bound based on Markov’s inequality for polynomials. This bound is independent of geometric “chain” constructions and isolates a limitation coming solely from polynomial degree constraints within the oblivious span model.
For fully adaptive first-order methods (allowed to use the entire oracle transcript), we will also show a stronger geometric $\Omega(1/T)$ min-iterate lower bound under a bounded scalarization gap by lifting the stationary-point lower bounds of Carmon–Duchi–Hinder–Sidford [7].
Theorem 4.7 (Polynomial-degree lower bound for oblivious span methods (last iterate)).
Fix $L,R>0$ and $T\in\mathbb{N}$. For any oblivious span first-order method run for $T$ steps, there exists an MOO instance with $L$-smooth convex objectives and $\mathrm{dist}(x^{(0)},\mathcal{P})=R$ such that the last iterate obeys
@@
By Lemma 4.2(c), there exists an $L$-smooth convex quadratic $g$ and an initialization $u^{(0)}$ such that for any oblivious span method run for $T$ steps,
$\|\nabla g(u^{(T)})\|\;\geq\;\frac{LR}{2\,(T+1)^{2}}\qquad\text{and}\qquad\mathrm{dist}\!\bigl(u^{(0)},\arg\min g\bigr)=R.$
Choose $u^{\star}\in\arg\min g$ to be a closest minimizer to $u^{(0)}$, so that $\|u^{(0)}-u^{\star}\|=\mathrm{dist}(u^{(0)},\arg\min g)=R$.
Apply Theorem 4.1 with $\gamma=L$ to lift $g$ into a non-degenerate MOO instance $(f_{1},\dots,f_{m})$. By Theorem 4.1(iv),
$\mathcal{P}=(\arg\min g)\times\mathrm{conv}\{a_{1},\dots,a_{m}\}.$
Pick $w^{(0)}\in\mathrm{conv}\{a_{1},\dots,a_{m}\}$ (e.g., $w^{(0)}=a_{1}$) and set $x^{(0)}=(u^{(0)},w^{(0)})\in V\oplus W$. Then
$\mathrm{dist}\!\bigl(x^{(0)},\mathcal{P}\bigr)=\mathrm{dist}\!\bigl(u^{(0)},\arg\min g\bigr)=R.$
Finally, Theorem 4.1(iii) gives $\mathcal{G}(x)\geq\|\nabla g(x_{V})\|$ for all $x$. Therefore,
$\mathcal{G}(x^{(T)})\;\geq\;\|\nabla g(x_{V}^{(T)})\|\;\geq\;\frac{LR}{2\,(T+1)^{2}},$
as claimed. ∎
Remark 4.8.
@@
Remark 4.9 (The polynomial-degree bound is loose).
Theorem 4.7 is a last-iterate bound derived from polynomial extremal inequalities and does not exploit geometric hard instances (e.g., zero-chains). As we show next, geometric stationary-point lower bounds imply a stronger min-iterate $\Omega(1/T)$ lower bound for adaptive methods in smooth convex MOO under a bounded scalarization gap.
We now lift the convex stationary-point lower bound of Carmon–Duchi–Hinder–Sidford [7]. The appropriate “value gap” assumption in MOO is expressed through a scalarization $f_{\lambda}$.
Our lifting construction (Theorem 4.1) produces a multiobjective instance $f_{1},\dots,f_{m}$ from a scalar function $g$ such that the Pareto stationarity gap dominates the scalar gradient norm, i.e., $\mathcal{G}(x)\geq\|\nabla g(x_{V})\|$ for every $x=(x_{V},x_{W})$. For lower bounds proved for restricted algorithm classes (e.g., oblivious one-step or (oblivious) span methods), this domination property is sufficient: the iterates must remain in prescribed linear spans / polynomial images of past first-order information, so the $V$-component of the iterates on the lifted instance inherits the same restriction, and the corresponding scalar lower bounds transfer directly via $\mathcal{G}(x)\geq\|\nabla g(x_{V})\|$.
In contrast, the lower bounds under a bounded initial value gap $\Delta$ (such as those of [7]) are minimax statements over all deterministic first-order methods. Here an additional reduction is required because a multiobjective first-order oracle returns strictly richer information than a scalar oracle: at each query $x$ it provides the full collection $\{(f_{i}(x),\nabla f_{i}(x))\}_{i=1}^{m}$, whereas in the scalar setting one observes only $(g(x_{V}),\nabla g(x_{V}))$. To legitimately invoke such scalar minimax lower bounds, we therefore use a simulation (transcript) argument: for the lifted family, any deterministic multiobjective algorithm can be simulated by a deterministic scalar first-order method on $g$ with the same number of oracle calls (since the multiobjective replies can be reconstructed from $(g(x_{V}),\nabla g(x_{V}))$ and the known quadratic $W$-terms). Consequently, if a multiobjective method were able to beat the scalar value-gap lower bound on the lifted instances, the simulated scalar method would also beat it, yielding a contradiction.
Theorem 4.10 (Geometric convex lower bound for adaptive methods under a bounded scalarization gap).
Fix $L>0$, $\Delta>0$, $T\in\mathbb{N}$, and $m\geq 2$. Consider the standard vector first-order oracle for MOO that, at any query $x\in\mathbb{R}^{d}$, returns the transcript item
$\mathcal{O}_{F}(x)\;:=\;\bigl((f_{i}(x),\nabla f_{i}(x))\bigr)_{i=1}^{m}.$
Let $\mathsf{Alg}$ be any deterministic first-order method that makes $T$ oracle calls to $\mathcal{O}_{F}$ and outputs iterates $x^{(0)},x^{(1)},\ldots,x^{(T)}$. Then there exists an $m$-objective MOO instance $F=(f_{1},\ldots,f_{m})$ with the following properties:
@@
    Each $f_{i}$ is convex and $L$-smooth, the objectives are distinct, and the Pareto set $\mathcal{P}$ is non-singleton (in fact, it contains a non-singleton polytope subset).
@@
    (Bounded scalarization gap.) For the scalarization $\lambda=e_{1}$,
    $\Delta_{e_{1}}(x^{(0)})\;:=\;f_{e_{1}}(x^{(0)})-\inf_{x}f_{e_{1}}(x)\;=\;f_{1}(x^{(0)})-\inf_{x}f_{1}(x)\;\leq\;\Delta.$
  3. (iii)
    (Min-iterate lower bound.) The iterates satisfy
    $\min_{0\leq t\leq T}\mathcal{G}(x^{(t)})\;\geq\;\frac{\sqrt{L\Delta}}{4(T+1)}.$
Equivalently, any deterministic first-order method needs $T=\Omega(\sqrt{L\Delta}/\varepsilon)$ oracle calls to guarantee $\min_{t\leq T}\mathcal{G}(x^{(t)})\leq\varepsilon$ on this class of instances.
@@
Oracle model as deterministic transcript maps. A deterministic first-order method $\mathsf{Alg}$ interacting with $\mathcal{O}_{F}$ can be formalized as a sequence of deterministic maps
$\Phi_{t}:\Big(\mathbb{R}^{d}\times(\mathbb{R}\times\mathbb{R}^{d})^{m}\Big)^{t}\to\mathbb{R}^{d},\qquad t\geq 0,$
such that, for $t=0,1,\ldots,T-1$,
$x^{(t+1)}=\Phi_{t}\Big(x^{(0)},\mathcal{O}_{F}(x^{(0)}),\ldots,x^{(t)},\mathcal{O}_{F}(x^{(t)})\Big).$
This is the standard deterministic first-order oracle model; $\mathsf{Alg}$ may use all past queried points and oracle replies.
Step 1: Choose a scalar hard instance from Carmon et al. Apply Lemma 4.3 to the deterministic first-order method that will be constructed below (it will make $T$ scalar first-order oracle calls). Thus there exists a convex $L$-smooth function $g:V\to\mathbb{R}$ (with $\dim(V)\geq 2T$) and an initialization $u^{(0)}\in V$ satisfying
$g(u^{(0)})-\inf_{u\in V}g(u)\leq\Delta$ (12)
and such that the iterates produced by that scalar method obey
$\min_{0\leq t\leq T}\|\nabla g(u^{(t)})\|\;\geq\;\frac{\sqrt{L\Delta}}{4(T+1)}.$ (13)
Step 2: Lift $g$ to a non-degenerate MOO instance. Let $W$ be an orthogonal complement of $V$ with $\dim(W)\geq m-1$ and pick affinely independent points $a_{1},\ldots,a_{m}\in W$. Define $F=(f_{1},\ldots,f_{m})$ on $\mathbb{R}^{d}=V\oplus W$ by the lifting construction (Theorem 4.1) with $\gamma=L$:
$f_{i}(u,w)\;:=\;g(u)+\frac{L}{2}\|w-a_{i}\|^{2},\qquad i=1,\ldots,m.$
By Theorem 4.1(i)–(iv), each $f_{i}$ is convex and $L$-smooth, the objectives are distinct, and the Pareto set is
$\mathcal{P}=(\arg\min g)\times\mathrm{conv}\{a_{1},\ldots,a_{m}\},$
hence non-singleton. In particular, for any $u^{\star}\in\arg\min g$, the slice $\{u^{\star}\}\times\mathrm{conv}\{a_{1},\dots,a_{m}\}\subseteq\mathcal{P}$ is a non-singleton polytope.
We choose the initialization
$x^{(0)}:=(u^{(0)},a_{1}).$
Step 3: Verify the bounded scalarization gap $\Delta_{e_{1}}(x^{(0)})\leq\Delta$. For $\lambda=e_{1}$, $f_{e_{1}}=f_{1}$ and
$f_{1}(u,w)=g(u)+\frac{L}{2}\|w-a_{1}\|^{2}.$
Hence $\inf_{(u,w)}f_{1}(u,w)=\inf_{u}g(u)$ (attained at $(u^{\star},a_{1})$), and therefore
$\Delta_{e_{1}}(x^{(0)})=f_{1}(u^{(0)},a_{1})-\inf_{(u,w)}f_{1}(u,w)=g(u^{(0)})-\inf_{u}g(u)\leq\Delta,$
by (12). This proves item (ii). Importantly, this does not make the MOO instance degenerate: for $i\neq 1$, the objectives differ because $a_{i}\neq a_{1}$. The role of $e_{1}$ is only to express a scalar value-gap condition needed to invoke Lemma 4.3.
Step 4: A transcript-level reduction from MOO to the scalar oracle for $g$. Let $\mathcal{O}_{g}$ be the scalar first-order oracle for $g$, returning $\mathcal{O}_{g}(u):=(g(u),\nabla g(u))$. We now define a deterministic scalar first-order method $\widetilde{\mathsf{Alg}}$ interacting with $\mathcal{O}_{g}$ that simulates the transcript of $\mathsf{Alg}$ interacting with $\mathcal{O}_{F}$ on the lifted instance.
Inductively suppose $\widetilde{\mathsf{Alg}}$ has constructed points $(u^{(0)},w^{(0)}),\ldots,(u^{(t)},w^{(t)})$ equal to the iterates of $\mathsf{Alg}$ on $F$. It queries $\mathcal{O}_{g}$ at $u^{(t)}$ and receives $(g(u^{(t)}),\nabla g(u^{(t)}))$. Using the known lift parameters $L$ and $(a_{i})_{i=1}^{m}$, it reconstructs the full vector-oracle reply at $x^{(t)}=(u^{(t)},w^{(t)})$ as
$\mathcal{O}_{F}(x^{(t)})=\Bigl(g(u^{(t)})+\tfrac{L}{2}\|w^{(t)}-a_{i}\|^{2},\ (\nabla g(u^{(t)}),\,L(w^{(t)}-a_{i}))\Bigr)_{i=1}^{m}.$
Since $\mathsf{Alg}$ is deterministic and its next iterate is given by the deterministic map $\Phi_{t}$ applied to the transcript, $\widetilde{\mathsf{Alg}}$ can compute
$x^{(t+1)}=\Phi_{t}\Big(x^{(0)},\mathcal{O}_{F}(x^{(0)}),\ldots,x^{(t)},\mathcal{O}_{F}(x^{(t)})\Big)$
and set $(u^{(t+1)},w^{(t+1)})=x^{(t+1)}$. Thus the sequence $(u^{(t)})_{t=0}^{T}$ generated by $\widetilde{\mathsf{Alg}}$ is exactly the $V$-projection of the iterates generated by $\mathsf{Alg}$ on the lifted instance. Moreover, $\widetilde{\mathsf{Alg}}$ makes exactly one scalar oracle call per iteration, hence $T$ calls total.
Therefore, the Carmon lower bound (13) applies to the simulated sequence $(u^{(t)})$.
Step 5: Transfer to the Pareto stationarity gap. By Theorem 4.1(iii), for all $t$,
$\mathcal{G}(x^{(t)})\;\geq\;\|\nabla g(u^{(t)})\|.$
Taking the minimum over $0\leq t\leq T$ and using (13) gives
$\min_{0\leq t\leq T}\mathcal{G}(x^{(t)})\;\geq\;\min_{0\leq t\leq T}\|\nabla g(u^{(t)})\|\;\geq\;\frac{\sqrt{L\Delta}}{4(T+1)}.$
This proves item (iii) and completes the proof. ∎
Theorem 4.10 is naturally stated in terms of a scalarization value gap $\Delta$. On our lifted family one can calibrate $\Delta$ and the Pareto distance scale $R$ simultaneously, yielding the $LR/T$ form.
Corollary 4.11 (From $\Delta$ to $R$ on the lifted family).
Fix $L>0$, $R>0$, $T\in\mathbb{N}$, and $m\geq 2$. There exists an MOO instance with $m$ distinct convex $L$-smooth objectives and an initialization $x^{(0)}$ such that:
  1. (i)
    $\mathrm{dist}(x^{(0)},\mathcal{P})=R$ and $\mathcal{P}$ is non-singleton (in particular, it contains a non-singleton polytope subset);
  2. (ii)
    for $\lambda=e_{1}$ the scalarization gap satisfies $\Delta_{e_{1}}(x^{(0)})\asymp LR^{2}$ (up to universal constants);
  3. (iii)
    for any deterministic first-order method run for $T$ steps,
    $\min_{0\leq t\leq T}\mathcal{G}(x^{(t)})\;\geq\;c_{0}\,\frac{LR}{T+1}$
    for a universal constant $c_{0}>0$.
Proof.
Let $(g_{0},u_{0}^{(0)})$ be the base hard instance from Theorem 4.10 for parameters $(L_{0},\Delta_{0})$, and denote $u_{0}^{\star}\in\arg\min g_{0}$ and $R_{0}:=\|u_{0}^{(0)}-u_{0}^{\star}\|>0$. For $a,s>0$, define the scaled function $g(u):=a\,g_{0}(su)$ and initialization $u^{(0)}:=u_{0}^{(0)}/s$. Then $\nabla g$ is $L=as^{2}L_{0}$-Lipschitz and
$g(u^{(0)})-\inf g=a\bigl(g_{0}(u_{0}^{(0)})-\inf g_{0}\bigr)=a\Delta_{0}=:\Delta.$
Moreover, since $\arg\min g=\{u_{0}^{\star}/s:u_{0}^{\star}\in\arg\min g_{0}\}$, we have
$\mathrm{dist}(u^{(0)},\arg\min g)=\frac{1}{s}\,\mathrm{dist}(u_{0}^{(0)},\arg\min g_{0})=\frac{R_{0}}{s}.$
Choose $s:=R_{0}/R$ and $a:=L/(s^{2}L_{0})$. Then $\nabla g$ is $L$-Lipschitz and
$\Delta=a\Delta_{0}=\frac{L\Delta_{0}}{s^{2}L_{0}}=\frac{\Delta_{0}}{L_{0}R_{0}^{2}}\,LR^{2}\asymp LR^{2},$
where the implied constants depend only on the fixed base instance.
Now lift $g$ using Theorem 4.1 with $m\geq 2$ distinct points $a_{1},\dots,a_{m}$ in $W$ and parameter $\gamma=L$, and set the initialization $x^{(0)}:=(u^{(0)},a_{1})$. By Theorem 4.1, the Pareto set is $\mathcal{P}=\arg\min g\times\mathrm{conv}\{a_{i}\}$, hence it is non-singleton, and since $(u^{\star},a_{1})\in\mathcal{P}$ for any $u^{\star}\in\arg\min g$ we have
$\mathrm{dist}(x^{(0)},\mathcal{P})=\mathrm{dist}(u^{(0)},\arg\min g)=R.$
Also, for $\lambda=e_{1}$ we have $f_{\lambda}=f_{1}(u,w)=g(u)+\frac{L}{2}\|w-a_{1}\|^{2}$, so $\Delta_{e_{1}}(x^{(0)})=g(u^{(0)})-\inf g=\Delta\asymp LR^{2}$.
Finally, applying Theorem 4.10 yields
$\min_{0\leq t\leq T}\mathcal{G}(x^{(t)})\geq\frac{\sqrt{L\Delta}}{4(T+1)}\asymp\frac{LR}{T+1},$
which proves the claim for a universal constant $c_{0}>0$. ∎
4.5 Nonconvex MOO with Lipschitz gradients
We now use the convexity-free nature of Theorem 4.1 to transfer classical nonconvex stationary-point lower bounds to Pareto stationarity. We emphasize that in nonconvex MOO, $\mathcal{G}(x)=0$ is a necessary stationarity condition (Lemma 3.6) but generally not sufficient for Pareto optimality.
Theorem 4.12 (Nonconvex MOO lower bound under $L$-Lipschitz gradients).
There exists a universal constant $c>0$ such that for any $L>0$, $\Delta>0$, $T\in\mathbb{N}$, and $m\geq 2$, the following holds. For any deterministic first-order method making $T$ vector-oracle calls, there exists an MOO instance with $m$ distinct objectives, each $L$-smooth (not necessarily convex), and an initialization $x^{(0)}$ such that:
  1. (i)
    For the scalarization $\lambda=e_{1}$, $\Delta_{e_{1}}(x^{(0)})\leq\Delta$.
  2. (ii)
    The iterates satisfy
    $\min_{0\leq t\leq T}\mathcal{G}(x^{(t)})\;\geq\;c\,\sqrt{\frac{L\Delta}{T+1}}.$
Equivalently, any deterministic first-order method requires $T=\Omega(L\Delta/\varepsilon^{2})$ oracle calls to guarantee $\min_{t\leq T}\mathcal{G}(x^{(t)})\leq\varepsilon$.
Proof.
Apply Lemma 4.4 to obtain an $L$-smooth (possibly nonconvex) scalar function $g:V\to\mathbb{R}$ with $g(u^{(0)})-\inf g\leq\Delta$ but
$\min_{0\leq t\leq T}\|\nabla g(u^{(t)})\|\geq c\sqrt{\frac{L\Delta}{T+1}}$
for any deterministic first-order method run for $T$ steps.
Lift $g$ to an MOO instance via Theorem 4.1 with $\gamma=L$ and initialization $x^{(0)}=(u^{(0)},w^{(0)})$ where $w^{(0)}\in\mathrm{conv}\{a_{i}\}$. As in the convex proof, $\Delta_{e_{1}}(x^{(0)})=g(u^{(0)})-\inf g\leq\Delta$. Moreover, by (7),
$\mathcal{G}(x^{(t)})\geq\|\nabla g(u^{(t)})\|\quad\forall t.$
Finally, the same simulation argument as in Theorem 4.10 shows that the $V$-coordinates of any MOO first-order method on the lifted instance correspond to a valid scalar first-order method on $g$, so the scalar lower bound transfers:
$\min_{0\leq t\leq T}\mathcal{G}(x^{(t)})\geq\min_{0\leq t\leq T}\|\nabla g(u^{(t)})\|\geq c\sqrt{\frac{L\Delta}{T+1}}.$